A three digit number is selected at random. What is the probability
that its unit digit is 2

Answer by sudhanshu_kmr(1152)


total no. of 3 digit number = 9*10*10 = 900
no. of 3 digit number having 2 at unit place = 9 * 10 * 1 = 90

probability = 90/900 = 1/10